### v1

# Subtlety of oscillation indices of oscillatory integrals of real analytic functions

Abstract.
For a locally defined real analytic function, we study the relation between the oscillation index of oscillatory integrals and the real log canonical threshold. The former is always negative, and its absolute value is greater than or equal to the latter. They coincide very often, but there are certain exceptional cases even in the Newton nondegenerate convenient homogeneous case where the number of variables is even and smaller than the degree.
This work was partially supported by National Research Foundation of Korea NRF-2023R1A2C1003390 and NRF-2022M3C1C8094326.
Introduction
Let $f$ be a real analytic function defined on a small open neighborhood $U$ of 0 in $X\,{:=}\,{\mathbb{R}}^{n}$ with f(0)= 0f(0)\,{=}\,0. For $C^{\infty}$ functions $\varphi$ on $U$ with compact supports, we can consider the oscillatory integrals
(1) $I(\tau,\varphi)\,{:=}\,\int_{X}e^{\sqrt{-1}\,\tau f(x)}\varphi(x){\rm d}{\mathbf{x}}\quad\quad\hbox{for}\,\,\,\tau\,{\in}\,{\mathbb{R}}_{>0},$
where ${\rm d}{\mathbf{x}}\,{:=}\,{\rm d}x_{1}{\wedge}\cdots{\wedge}{\rm d}x_{n}$ with ${\mathbf{x}}\,{=}\,(x_{1},\dots,x_{n})$ the coordinate system of ${\mathbb{R}}^{n}$. As a corollary of Hironaka’s resolution of singularities [Hi 64], we have asymptotic expansions  for $\tau\to+\infty$ :
(2) $I(\tau,\varphi)\sim\sum_{\alpha<0}\sum_{k=0}^{n-1}C_{\alpha,k}(\varphi)\tau^{\alpha}(\log\tau)^{k}\quad\quad\bigl(C_{\alpha,k}(\varphi)\,{\in}\,{\mathbb{C}}\bigr),$
see [Ma 74], [Va 76], [AGV 88], [CKN 13]. The oscillation index  $\beta(f)$ of $f$ at $0\,{\in}\,X$ is by definition the maximal number $\beta$ such that for any neighborhood $V$ of 0 in $U$, there is a $C^{\infty}$ function $\varphi$ compactly supported in $V$ with $C_{\beta,k}(\varphi)\,{\neq}\,0$ for some $k$.
On the other hand, we have the real log canonical threshold  ${\rm rlct}(f)$ at 0 for a real analytic function $f$ defined on $U$, which can be determined also by using a resolution of singularities, see [Sa 07]. Indeed, let $\pi\,{:}\,\widetilde{U}\to U$ be a resolution of singularities for $f$, which is obtained by repeating smooth center blowups, and such that $\pi^{*}\!f$ and its Jacobian determinant are monomials up to multiplication by unit for some local coordinates at each point of $\pi^{-1}f^{-1}(0)$, see [Hi 64]. Let $D_{j}$ ($j\,{\in}\,J$) be the irreducible components of $D\,{:=}\,\pi^{-1}f^{-1}(0)$, which we may assume to be smooth. Let $m_{j},k_{j}$ be the multiplicities of $\pi^{*}\!f$ and $\pi^{*}{\rm d}{\mathbf{x}}$ at general points of $D_{j}$. Set
(3) $\gamma(f)\,{:=}\,\min_{j\,{\in}\,J}\bigl\{\tfrac{k_{j}{+}1}{m_{j}}\bigr\}.$
Note that $\gamma(f)\,{\leqslant}\,1$ if $f^{-1}(0)$ is (n1)(n{-}1)-dimensional. We have the following
Theorem 1 ([Sa 07]).
There is an equality
(4) ${\rm rlct}(f)=\gamma(f).$
Note that the real log canonical threshold can be greater than 1 and also than the log canonical threshold ${\rm lct}(f_{{\mathbb{C}}})$ in the case $\dim f^{-1}(0)\,{<}\,n{-}1$, where $f_{{\mathbb{C}}}$ is the complexification of $f$ obtained by the inclusion ${\mathbb{R}}\{{\mathbf{x}}\}\hookrightarrow{\mathbb{C}}\{{\mathbf{x}}\}$.
On the other hand, we have the following.
Theorem 2 ([Va 76]).
There is an inequality
(5) $\beta(f)\,{\leqslant}\,{-}\gamma(f).$
Note that the normal crossing case is a very special case of a Newton nondegenerate case, see also [CKN 13]. It seems to be believed by some people that the equality holds in (5) at least in the Newton nondegenerate case with $\gamma(f)\,{<}\,1$.
In this note we show the following.
Theorem 3.
Assume $f$ is Newton $\mathbb{R}$}-nondegenerate, convenient, homogeneous, $n$ is even, and either $d\,{:=}\,\deg f\,{>}\,n$ or $f^{-1}(0)\,{=}\,\{0\}$ ((with $d$ even)). Then we have $\gamma(f)\,{=}\,\tfrac{n}{d}$ together with the strict inequality
(6) $\beta(f)\,{<}\,{-}\gamma(f).$
Note that if $f$ is homogeneous and $d$ is odd, then $\dim f^{-1}(0)\,{=}\,n{-}1$, see Remark  3.1 below. (Theorem  3 does not seem quite compatible with some well-known assertion in the literature, see for instance [CKN 13].)
In Section 1, we review some basics of (real) jumping numbers and (real) log canonical thresholds. In Section 2 we recall some theorem from the theory of oscillatory integrals. In Section 3 we prove Theorem 3 using the assertions in the previous section.
1. Log canonical thresholds
In this section we review some basics of (real) jumping numbers and (real) log canonical thresholds.
Let $f\,{\in}\,{\mathbb{R}}\{{\mathbf{x}}\}\,{=}\,{\mathbb{R}}\{x_{1},\dots,x_{n}\}$. This real analytic function is defined on a small open subset $U$ of $X\,{:=}\,{\mathbb{R}}^{n}$. We have the real multiplier ideals  ${\mathcal{J}}(U,f^{\alpha})\,{\subset}\,{\mathcal{O}}_{U}$ ($\alpha\,{\in}\,{\mathbb{R}}_{\geqslant 0}$) defined by the local integrability  of $|g|/|f|^{\alpha}$ for $g\in{\mathcal{O}}_{U}$. (Here coherence of ${\mathcal{J}}(U,f^{\alpha})$ is unclear.) These are weakly decreasing for $\alpha$, and coincide with ${\mathcal{O}}_{U}$ for $0\,{\leqslant}\,\alpha\,{\ll}\,1$. It is known that there is an increasing sequence of positive rational numbers $\alpha_{k}$ ($k\,{\in}\,{\mathbb{Z}}_{>0}$), called the real jumping numbers  of $f$, such that
(1.1) ${\mathcal{J}}(U,f^{\beta})={\mathcal{J}}(U,f^{\alpha_{k}})\neq{\mathcal{J}}(U,f^{\alpha_{k+1}})\quad\hbox{if}\,\,\,\beta\,{\in}\,[\alpha_{k},\alpha_{k+1}),\,k\,{\in}\,{\mathbb{N}}$,
where $\alpha_{0}\,{=}\,0$. The minimal real jumping number is called the real log canonical threshold, and is denoted by ${\rm rlct}(f)$, see [Sa 07] (where ${\mathbb{R}}\{\!\{x\}\!\}$ means ${\mathbb{R}}\{{\mathbf{x}}\}$).
Remark 1.1.
We do not  necessarily have
(1.2) $f{\mathcal{J}}(U,f^{\alpha})={\mathcal{J}}(U,f^{\alpha+1})\quad\hbox{for}\,\,\,\alpha\,{>}\,0,$
in the exceptional type case. Here we say that $f$ is exceptional type if ${\rm codim}_{U}\,D\,{>}\,1$ with $D\,{:=}\,f^{-1}(0)\,{\subset}\,U$, and ordinary type  otherwise.
Remark 1.2.
The above equality (1.2) always holds for the complexification $f_{{\mathbb{C}}}$ of $f$, which is defined on a sufficiently small open neighborhood $U_{{\mathbb{C}}}$ of 0 in ${\mathbb{C}}^{n}$ with $U_{{\mathbb{C}}}\cap{\mathbb{R}}^{n}\,{=}\,U$ using the inclusion ${\mathbb{R}}\{{\mathbf{x}}\}\hookrightarrow{\mathbb{C}}\{{\mathbf{x}}\}$. The multiplier ideals ${\mathcal{J}}(U_{{\mathbb{C}}},f_{{\mathbb{C}}}^{\,\alpha})$ can be defined by using the local integrability of $|g|^{2}/|f_{{\mathbb{C}}}|^{2\alpha}$ for $g\,{\in}\,{\mathcal{O}}_{U_{{\mathbb{C}}}}$, and they are closely related to the $V$-filtration of Kashiwara and Malgrange indexed by  $\mathbb{Q}$}, see for instance [BuSa 05].
Remark 1.3.
The minimal jumping number is called the log canonical threshold, which is denoted by ${\rm lct}(f_{{\mathbb{C}}})$. This coincides with the minimal spectral number in the isolated singularity case if the latter is at most 1, see for instance [JKSY 22]. We have an inequality
(1.3) ${\rm lct}(f_{{\mathbb{C}}})\leqslant{\rm rlct}(f),$
where the equality does not necessarily hold, see [Sa 07] (and also Remark  3.3 below). For the relation to the maximal root of Bernstein-Sato polynomial up to sign, see [Ko 97] (and [JKSY 22, Remark 1.2 (iii)]).
2. Oscillatory integrals
In this section we recall some theorem from the theory of oscillatory integrals.
For $f,\varphi$ as in the introduction, let $\Gamma_{\!+}(f),\Gamma_{\!+}(\varphi)$ be their Newton polytopes  defined by using their Taylor expansions with coordinates $x_{1},\dots,x_{n}$ fixed; for instance $\Gamma_{\!+}(f)$ is the convex hull of the union of $\nu\thinspace{+}\thinspace{\mathbb{R}}_{\geqslant 0}^{\,n}\,{\subset}\,{\mathbb{R}}_{\geqslant 0}^{\,n}$ for $\nu\,{\in}\,{\mathbb{N}}^{n}$ with $a_{\nu}\,{\neq}\,0$, where the $a_{\nu}$ are the coefficients of the Taylor expansion $\hbox{$\sum$}_{\nu\in{\mathbb{N}}^{n}}\,a_{\nu}{\mathbf{x}}^{\nu}$ of $f$ with ${\mathbf{x}}^{\nu}\,{:=}\,\hbox{$\prod$}_{i=1}^{n}\,x_{i}^{\nu_{i}}$.
A real analytic function $f$ is called Newton $\mathbb{R}$}-nondegenerate  if for any compact face $\sigma$ of $\Gamma_{\!+}(f)$, the intersection of $\{\partial_{x_{i}}f_{\sigma}\,{=}\,0\}\,{\subset}\,{\mathbb{R}}^{n}$ for $i\,{\in}\,[1,n]$ is contained in $\{x_{1}\cdots x_{n}\,{=}\,0\}$, where $f_{\sigma}\,{:=}\,\hbox{$\sum$}_{\nu\in\sigma}\,a_{\nu}{\mathbf{x}}^{\nu}$ is the partial sum of the Taylor expansion of $f$ over $\sigma\cap{\mathbb{N}}^{n}$. We define $\mathbb{C}$}-nondegeneracy considering the zero-locus in ${\mathbb{C}}^{n}$ instead of ${\mathbb{R}}^{n}$. We say that $f$ is convenient  if the Newton polytope $\Gamma_{\!+}(f)$ intersects every coordinate axis of ${\mathbb{R}}^{n}$, see [Va 76], [CKN 13].
Assume $\Gamma_{\!+}(f),\Gamma_{\!+}(\varphi)$ are non-empty, or equivalently, their Taylor expansions are nonzero. Set
$d(f,\varphi):=\min\bigl\{d\,{\in}\,{\mathbb{R}}_{>0}\mid d\,{\cdot}\,(\Gamma_{\!+}(\varphi){+}{\bf 1})\,{\subset}\,\Gamma_{\!+}(f)\bigr\},$
where ${\bf 1}\,{:=}\,(1,\dots,1)\,{\in}\,{\mathbb{R}}^{n}$. Set $({\mathbb{R}}_{\geqslant 0}^{\,n})^{\geqslant r}\,{:=}\,\{\nu\,{\in}\,{\mathbb{R}}_{\geqslant 0}^{\,n}\mid|\nu|\,{\geqslant}\,r\}$ with $|\nu|\,{:=}\,\hbox{$\sum$}_{i=1}^{n}\nu_{i}$.
If $f$ is convenient so that
$\Gamma_{\!+}(f)\supset({\mathbb{R}}_{\geqslant 0}^{\,n})^{\geqslant r},\quad\Gamma_{\!+}(\varphi)\subset({\mathbb{R}}_{\geqslant 0}^{\,n})^{\geqslant r^{\prime}},$
for some $r,r^{\prime}\,{\in}\,{\mathbb{R}}_{>0}$, then we have
(2.1) $d(f,\varphi)\leqslant\frac{r}{r^{\prime}+n}.$
Note that r=dr\,{=}\,d if $f$ is a convenient homogeneous polynomial of degree $d$.
Let $\beta(f,\varphi)$ be the maximal number $\beta$ such that $C_{\beta,k}(\varphi)\,{\neq}\,0$ for some $k$ in (1), see [CKN 13]. We have the following.
Theorem 2.1 ([CKN 13, Theorem 2.2]).
Assume $f$ is convenient, Newton $\mathbb{R}$}-nondegenerate, and $\Gamma_{\!+}(\varphi)\,{\neq}\,\emptyset$. Let $r,r^{\prime}$ be as in (2.1). Then we have the inequalities
(2.2) $\beta(f,\varphi)\leqslant-1/d(f,\varphi)\leqslant-\frac{r^{\prime}+n}{r}.$
We say that a compactly supported $C^{\infty}$ function $\chi$ on $U$ is a cutoff function  if the values of $\chi$ are nonnegative and coincide with 1 on a neighborhood of 0 in $U$. (The support of $\varphi$ is the closure of the nonzero locus in $U$.) Theorem  2.1 implies the following.
Corollary 2.1.
Assume $f$ is Newton $\mathbb{R}$}-nondegenerate and convenient. Then the asymptotic expansion of $I(\tau,{\mathbf{x}}^{\nu}\chi)$ is independent of the choice of a cutoff function $\chi$ for any $\nu\,{\in}\,{\mathbb{N}}^{n}$.
Proof.
Let $\varphi_{k}$ be compactly supported $C^{\infty}$ functions on $U$ such that $({\mathbb{R}}_{\geqslant 0}^{\,n})^{\geqslant k}\,{\supset}\,\Gamma_{\!+}(\varphi_{k})\,{\neq}\,\emptyset$ for $k\,{\gg}\,0$. This corollary can be verified by applying Theorem  2.1 to the difference between ${\mathbf{x}}^{\nu}\chi\pm\varphi_{k}$ and ${\mathbf{x}}^{\nu}\chi$, and also to the one between ${\mathbf{x}}^{\nu}\chi\thinspace{+}\thinspace\varphi_{k}$ and ${\mathbf{x}}^{\nu}\chi^{\prime}\thinspace{-}\thinspace\varphi_{k}$ for two cutoff functions $\chi$ and $\chi^{\prime}$, for any $k\,{\gg}\,0$. ∎
Corollary 2.2.
Under the hypothesis of Corollary  2.1, the asymptotic expansion of $I(\tau,\varphi)$ depends only on the Taylor expansion $\hbox{$\sum$}_{\nu\in{\mathbb{N}}^{n}}\,c_{\nu}{\mathbf{x}}^{\nu}$ of $\varphi$.
Proof.
This follows from Corollary  2.1 applying Theorem  2.1 to $\varphi-\hbox{$\sum$}_{|\nu|\leqslant k}\,c_{\nu}{\mathbf{x}}^{\nu}\chi$ for any $k\gg 0$. (Note that the Taylor expansion of $\varphi$ is not necessarily a convergent power series.) ∎
3. Proof of the main theorem
In this section we prove Theorem 3 using the assertions in the previous section.
Let $\pi\,{:}\,\widetilde{X}\,{\to}\,X$ be the blowup at $p\,{\in}\,X$. The embedded resolution of singularities is given by this. Set $E\,{:=}\,\pi^{-1}(0)$. For $i\,{\in}\,[1,n]$, there is a coordinate system ${\mathbf{y}}^{(i)}\,{=}\,(y^{(i)}_{1},\dots,y^{(i)}_{n})$ of the complement $V_{i}$ of the proper transform of $x_{i}^{-1}(0)$ in $\widetilde{X}$ such that
(3.1) $y^{(i)}_{i}\,{=}\,\pi^{*}x_{i},\quad y^{(i)}_{j}\,{=}\,\pi^{*}(x_{j}/x_{i})\,\,\,(j\,{\neq}\,i),$
$\pi^{*}\!f\,{=}\,(y^{(i)}_{i})^{d}\thinspace h_{i}(\widehat{{\mathbf{y}}}^{(i)}),\quad\pi^{*}{\rm d}{\mathbf{x}}\,{=}\,(y^{(i)}_{i})^{n-1}\thinspace{\rm d}{\mathbf{y}}^{(i)},$
with $h_{i}(\widehat{{\mathbf{y}}}^{(i)})\,{:=}\,f|_{x_{i}=1}$. Here $x_{j}$ for $j\,{\neq}\,i$ is identified with $y^{(i)}_{j}$ in the definition of $h_{i}(\widehat{{\mathbf{y}}}^{(i)})$, and $\widehat{{\mathbf{y}}}^{(i)}\,{:=}\,(y^{(i)}_{1},\dots,y^{(i)}_{i-1},y^{(i)}_{i+1},\dots,y^{(i)}_{n})$ for $i\,{\in}\,[2,n{-}1]$ (similarly for i= 1i\,{=}\,1 or $n$). We then get
(3.2) ${\rm rlct}(f)\,{=}\,\tfrac{n}{d},$
by Theorem  1.
We have the decomposition
(3.3) $V_{i}=E_{i}\times{\mathbb{R}},$
in a compatible way with the coordinates $\widehat{{\mathbf{y}}}^{(i)}$ and $y^{(i)}_{i}$, where $E_{i}$ is the complement in $E$ of the intersection of $E$ with the proper transform of $x_{i}^{-1}(0)$. For $i\,{\in}\,[1,n]$, set
$U^{\prime}_{i}\,{:=}\,\bigl\{|y^{(i)}_{j}|\,{<}\,1\thinspace{+}\thinspace\varepsilon\,\,\,(j\,{\neq}\,i)\bigr\}\subset E_{i},$
[…]
$\widetilde{\chi}\,{:=}\,\hbox{$\sum$}_{i=1}^{n}\,\eta(y^{(i)}_{i})\thinspace\theta_{i}(\widehat{{\mathbf{y}}}^{(i)}).$
The restriction of $\widetilde{\chi}$ to a sufficiently small neighborhood of $E$ is equal to the constant function with value 1. So there is a cutoff function $\chi^{\prime}$ on $X$ with $\pi^{*}\chi^{\prime}\,{=}\,\widetilde{\chi}$. Applying Theorem  2.1 to $\varphi\thinspace{-}\thinspace\varphi(0)\chi^{\prime}$ (and using a remark after (2.1)), the assertion is reduced to
(3.4) $\beta(f,\chi^{\prime})<-\tfrac{n}{d}.$
In the $d$ even case it is well sufficient to show for any $i\,{\in}\,[1,n]$ the vanishing
(3.5) $\int_{E_{i}}\!\int_{{\mathbb{R}}}e^{\sqrt{-1}\,\tau y_{i}^{d}h_{i}(\widehat{{\mathbf{y}}})}y_{i}^{n-1}\thinspace\eta(y_{i})\thinspace\theta_{i}(\widehat{{\mathbf{y}}})\thinspace{\rm d}y_{i}\thinspace{\rm d}\widehat{{\mathbf{y}}}=0,$
by the definition of $\chi^{\prime}$ (since $f$ is homogeneous). Here (i) is omitted to simplify the notation. This vanishing is however trivial using the double integral and calculating the integral over $\mathbb{R}$} first, since $d,n$ are even, and $\eta$ is an even function (and the integration of an odd function with compact support on $\mathbb{R}$} vanishes). Note that this is well compatible with [CKN 13, Section 7.1, (ii)].
In the $d$ odd case we can apply the above argument to the real part of $I(\tau,\chi^{\prime})$, since we have the identity
$e^{\sqrt{-1}\thinspace\theta}\,{=}\,\cos\theta\thinspace{+}\thinspace\sqrt{-1}\thinspace\sin\theta\quad(\theta\,{\in}\,{\mathbb{R}}),$
with $\cos\theta$, $\sin\theta$ respectively even  and odd  functions of $\theta$. We can show the vanishing of the first leading terms with respect to $\alpha$ of the asymptotic expansion of the imaginary part of $I(\tau,\chi^{\prime})$ (see (1)–(2)) using the involution $\iota$ of $X$ such that $\iota^{*}x_{i}\,{=}\,{-}x_{i}$ ($\forall\,i\,{\in}\,[1,n]$). Indeed, we have $\iota^{*}f\,{=}\,{-}f$ and we may assume $\iota^{*}\chi^{\prime}\,{=}\,\chi^{\prime}$ replacing $\chi^{\prime}$ with $(\chi^{\prime}\thinspace{+}\thinspace\iota^{*}\chi^{\prime})/2$ and applying Corollary  2.1. (Here $\iota^{*}{\rm d}{\mathbf{x}}\,{=}\,(-1)^{n}{\rm d}{\mathbf{x}}$, but we do not have to use the condition that $n$ is even, since an orientation  is needed to define the integration of a form of highest degree. Similarly a symmetry of $f$ does not imply the vanishing of an oscillatory integral using the action of a permutation of two coordinates.)
Remark 3.1.
One could get that $\beta(f)\,{=}\,{-}\tfrac{n{+}1}{d}$ under the hypotheses of Theorem  3 using [CKN 13, Section 7.1, (i) and (iii)]. The arguments as above could imply that the equality holds in (5) if $f$ is as in Theorem  3 with $f^{-1}(0)\,{=}\,\{0\}$, $d$ is even, $n$ is odd, and $\tfrac{n}{d}\,{<}\,1$, using [CKN 13, Section 7.1, (i)]. Here the complement of $f^{-1}(0)$ cannot be connected if $d$ is odd (using $\iota$). Note also that the values of the functions $h_{i}(\widehat{{\mathbf{y}}}^{(i)})$ for $\widehat{{\mathbf{y}}}^{(i)}\,{\in}\,\overline{U^{\prime}_{i}}$ are bounded.
Remark 3.2.
The above argument could be extended to the case $f$ is convenient, Newton $\mathbb{R}$}-nondegenerate, and nonnegative-valued (without assuming that $f$ is homogeneous). Let $\sigma_{j}$ ($j\,{\in}\,J_{0}$) be the (n1)(n{-}1)-dimensional faces of $\Gamma_{\!+}(f)$ containing the intersection of $\partial\Gamma_{\!+}(f)$ with the diagonal spanned by ${\bf 1}\,{\in}\,{\mathbb{R}}^{n}$. Let $d_{j}$ be the LCM of the denominators of the coefficients of the linear function $\ell_{j}$ on ${\mathbb{R}}^{n}$ such that $\sigma_{j}\,{\subset}\,\ell_{j}^{-1}(1)$. Set $r_{j}\,{=}\,d_{j}\ell_{j}({\bf 1})$. (Note that the pullback of a monomial ${\mathbf{x}}^{\nu}$ with $d_{j}\ell_{j}(\nu)\,{=}\,k$ has zero of order $k$ at general point of a divisor corresponding to $\sigma_{j}$, and the pullback of a logarithmic form ${\mathbf{x}}^{-{\bf 1}}{\rm d}{\mathbf{x}}$ is a logarithmic form.) Assuming $1\,{>}\,r_{j}/d_{j}\,{=}\,\ell_{j}({\bf 1})$ (which is independent of $j\,{\in}\,J_{0}$), the equality could hold in (5) with $\gamma(f)\,{=}\,r_{j}/d_{j}$ if $d_{j}$ is even and $r_{j}$ is odd for some $j\,{\in}\,J_{0}$. (Without the assumption on the non-negativity of $f$, it does not seem easy to show the non-cancellation  among the integrals associated with different faces $\sigma_{j}$ or between the integrals over $\{h_{\sigma}\,{>}\,0\}$ and $\{h_{\sigma}\,{<}\,0\}$.) Here one may have a problem about the commutativity of asymptotic expansion and integration.
Remark 3.3.
There is a big difference between Newton nondegeneracy over $\mathbb{R}$} and $\mathbb{C}$}; for instance in the case $f$ is homogeneous, convenient, and $f^{-1}(0)\,{=}\,\{0\}$, the $m$ th power $f^{m}$ is Newton $\mathbb{R}$}-nondegenerate for any positive integer $m$ if $f$ is. There is also an example like $f\,{=}\,(x^{2}\thinspace{+}\thinspace y^{2}\thinspace{+}\thinspace z^{2})^{2}\thinspace{+}\thinspace x^{6}\thinspace{+}\thinspace y^{6}\thinspace{+}\thinspace z^{6}$, which is Newton nondegenerate over $\mathbb{R}$}, but not over $\mathbb{C}$}. Here ${\rm rlct}(f)\,{=}\,\tfrac{3}{4}$, ${\rm lct}(f_{{\mathbb{C}}})\,{=}\,\tfrac{2}{3}$ (using Remark  1.3), and $\tfrac{3}{4}$ is not  a spectral number of $f_{{\mathbb{C}}}$ according to Singular [DGPS 23] (where the spectral numbers are shifted by 1-1).